OpenStudy (anonymous):

Write the equation of the ellipse 4x2 + 9y2 – 16x – 126y + 421 = 0 in standard form.

OpenStudy (mertsj):

1. Move the 421 to the right side. 2. Group the x^2 term and the x term together and factor out the 4. 3. Group the y^2 term and the y term together and factor out the 9 4. Complete the square twice. 5. Factor

OpenStudy (anonymous):

Thank you! xoxo

OpenStudy (mertsj):

yw

OpenStudy (mertsj):

Is it working out for you?

OpenStudy (anonymous):

I got stuck on the factoring part.

OpenStudy (mertsj):

Perhaps I can help.

OpenStudy (mertsj):

\[4(x^2-4x+4)+9(y^2-14y+49)=-421+16+441\]

OpenStudy (mertsj):

\[4(x+2)^2+9(y-7)^2=36\]

OpenStudy (mertsj):

Whoops. Make that x-2

OpenStudy (anonymous):

Okay, that makes sense. I've always been bad at factoring. Thank you/

OpenStudy (mertsj):

\[\frac{(x-2)^2}{9}+\frac{(y-7)^2}{4}=1\]